OpenStudy (anonymous):

List and describe each of the parts of the ideal gas law. Be sure to include the units for each part.

OpenStudy (abb0t):

The most basic one: PV = nRT where: p = pressure (atm) V = volume (usually in L) n = mole R = gas constant (u can find the units in ur book or online) T = temperature [K] what this says is that the internal energy depends on only a temperture and the number of particles; it oesn't depend on the volume occupied by the gas. This is the most simple form. Since gases are composed of molecules in random motion. if the molecules are perfectly elastic and the molecules DON'T inteeract w/ each other then the gas is said to be ideal. And by the equation, temperature is related to speed and related to the average kinetic energy of any molecule in the system. but remember, that no real gas EXACTLY satisfies this equation. life is much more complicated than that, in which you have an extended form of the ideal gas law, this takes into account intermolecular interaction, hence molecules are now attracted to each other, now you have: \(\left( \frac{P+an^2 }{ V^2 } \right)\left( V-nb \right)= nRT\), in this case however, a and b are constants which accounts for hte interaction forces between each other.

OpenStudy (anonymous):

Wow thank you!